# Object-Attribute-Relation Representation based Video Semantic Communication

### v1

@@ I Introduction
Research into explicit semantic representation and transmission of videos is still limited. Studies such as [16, 17, 18, 19, 20] have explored using semantic segmentation maps and scene graphs for semantic representations, implementing corresponding image and video generation to preserve the videos’ semantic content. However, these studies did not deeply engage with compression perspectives or data volume reduction. In [21, 22, 8], semantic segmentation maps served as explicit semantic representations for image compression, enhancing semantic fidelity by varying bit-rates across different categories. These efforts treated semantic information as supplementary to traditional image compression, relying still on transmitting numerous pixel and residual features for quality enhancements, which hindered significant bit-rate reductions. [23] took a different approach by encoding and transmitting only semantic segmentation maps and reconstructing images using a diffusion model, catering to extreme channel conditions. Similarly, [11] used sketch representations and generative models. In video conferencing, efficient transmission was achieved using facial keypoints[24]. Nevertheless, these approaches primarily focused on the semantic representation and coding of images or simpler video formats. The challenge remains to explore these techniques for large-scale, time-continuous, and object-consistent video representation to fully address the complexities of large scene videos.
@@
To address the above issues, this paper proposes a novel video transmission system leveraging object-attribute-relation (OAR) representation model of semantic information, as shown in Fig. 1 (b). Specifically, object-attribute-relation (OAR) model characterizes scene semantics by capturing key objects and their attributes and relations of video frames with regards to downstream application requirements. Compared with object edges and semantic segmentation maps widely used in previous studies, OAR captures more structured semantic correlations, which are represented by connected graphs; compared with scene graph representation, OAR pays more attention to the enrichment of attributes, which enables better reconstruction of videos. In addition, this paper implements a JSCC pipeline for video transmission by exploiting the advantages of compact and downstream tasks related OAR data. Specifically, this paper implements OAR-assisted JSCC for key frames so that the model can pay more attention to the key objects in the scene. Experimental validation is performed on traffic surveillance videos, and the results indicate that the proposed system can still maintain better foreground reconstruction quality under low bandwidth and SNR conditions.
@@
    Experiments on a traffic surveillance dataset show that the proposed method improves the perceptual quality of reconstructed videos significantly compared to conventional methods such as H.265 and prior work in semantic communication, namely DVST. Specifically, at a CBR of 1/300, the proposed method exhibits an average reduction in LPIPS loss of 0.037 when compared to DVST. Additionally, it facilitates improved performance in object detection tasks, achieving an average increase in mean average precision (mAP) of 0.16 compared to DVST.
@@ II-A Object-Attribute-Relation Formulation and OAR Based Video Communication Framework
An Object-Attribute-relation (OAR) representation is a structured abstraction of the scene semantics in the form of a graph. For image $\bf I$, the OAR is expressed as a tuple of objects, attributes and relations: ${\bf I}\stackrel{{\scriptstyle\mathcal{E}_{OAR}}}{{\longrightarrow}}\left(\mathcal{O},\mathcal{A},\mathcal{R}\right)$, where $\mathcal{E}_{OAR}$ denotes the mapping from the image to OAR. For video frames, the OAR is represented as a temporal sequence, $[\left(\mathcal{O}_{1},\mathcal{A}_{1},\mathcal{R}_{1}\right),\cdots,\left(\mathcal{O}_{T},\mathcal{A}_{T},\mathcal{R}_{T}\right)]$.
As shown in the formula, OAR contains three components: ${\mathcal{O}}=[o_{1},\cdots,o_{N}]$, ${\mathcal{A}}=[{\bf a}_{1},\cdots,{\bf a}_{N}]$, and =${\mathcal{R}}=\{r_{1},\cdots,r_{K}\}$, where $N,K$ denote the numbers of objects and relations, respectively. ${\mathcal{O}}$ is the object set with $o_{n}$ denoting a unique ID for each object. ${\mathcal{A}}$ is the attribute set, with ${\bf a}_{n}=\left(a^{(n)}_{1},a^{(n)}_{2},\cdots,a^{(n)}_{A}\right)={\mathcal{E}_{\text{A}}}({\bf I},o_{n})$ denoting $A$ attributes of object $o_{n}$. The relation is denoted by =${\mathcal{R}}=\left\{{\bf r}_{1},\cdots{\bf r}_{R}\right\}=\left\{(o_{\text{s},1},o_{\text{o},1},r_{1}),\cdots,(o_{\text{s},K},o_{\text{o},K},r_{K})\right\}$, $r_{k}=r_{o_{\text{s},k}\rightarrow o_{\text{s},k}}$ representing the relation with subject $o_{\text{s},k}$ and object $o_{\text{o},k}$. An OAR is characterized by a directed graph with objects as nodes and relations as edges. The relations and attributes (e.g., category, color, etc.) take values from a deterministic set, i.e.,
@@
OAR is highly interpretable and has a small data volume because only semantics of key objects are characterized. It is especially conducive to task-oriented applications with steady backgrounds, such as surveillance videos. Additionally, OAR provides scalability and can be adapted to more complex scenarios through the extension of attributes and relations. Based on the above formulation, this paper proposes an OAR-based video transmission framework, as presented in Fig. 2.
In the proposed framework, $T$ consecutive frames are encapsulated as one group-of-pictures (GoP). The first frame of each GoP (reference frame or key frame) is intra-coded and transmitted, providing contour and texture information about the background and key objects. OAR sequences are extracted for subsequent frames and transmitted. The receiver carries out generative reconstruction based on the reference frame incorporated with OAR sequences. Without loss of generality, each GOP is independently processed. Reference frames are denoted by ${\bf I}_{1}$, while subsequent $T-1$ frames are denoted by ${\bf I}_{t},t=2,\cdots,T$. For objects contained in reference frames, faithful reconstruction is realized based on the motion information in OAR. Newly appeared objects are generated based on the prior knowledge of the trained models, resulting in possible bias in details such as colors and textures. This paper focuses on task-oriented video transmission, so deviation in detailed textures are acceptable under the premise of semantic fidelity.
II-B OAR Formulation and Extraction
OAR representation is the critical component of the proposed video transmission system, providing semantics for low bit-rate representation and enabling efficient video transmission. Considering a surveillance video transmission system oriented towards tasks such as object detection and anomaly monitoring. Vehicles in the foreground are treated as the objects in OAR. Attributes consist of the position ${\bf p}=(x,y)$, size ${\bf s}=(w,h)$, angle $\theta$, and category $c$, denoted by ${\bf a}=(a_{1},a_{2},a_{3},a_{4})=({\bf p},{\bf s},\theta,c)$. Relations involve mainly “occlusions” between foreground objects and occlusions of the foreground by the background. The trivial relation of objects “in” the background is also included. Consequently, for an image of size $W\times H$, the attributes and relations take the value space of:
@@ II-B OAR Formulation and Extraction
To realize automated OAR extraction, multi-object tracking (MOT) is utilized, followed by several attribute recognition models. The overall process mainly consists of three parts: multi-object tracking, attribute recognition, and relation identification.
First, multi-object tracking. Object detection is first conducted via YOLOv9[28] to obtain the categories $c_{n}^{(t)}$ and bounding boxes ${\bf b}_{n}^{(t)}=(x_{n}^{(t)},y_{n}^{(t)},w_{n}^{(t)},h_{n}^{(t)})$, where $n=1,\cdots,N_{t}$. To match the identical objects in consecutive frames, a unique ID $o_{n}^{(t)}$ is assigned to each object via SORT [29]. IDs for all objects constitute the object component ${\mathcal{O}_{t}}=[o_{1}^{(t)},\cdots,o_{N_{t}}^{(t)}]$ in OAR.
@@
Third, relation identification. The occlusion relation is determined among object pairs with intersecting bounding boxes by comparing the appearance of the overlapping regions with the respective object features. Besides, occlusions between objects and background elements are also considered. Regions located at the spatial forefront are manually labeled. Objects overlapping with the labeled area are considered to be occluded by the background. Incorporating the identified “occlusion” and the trivial “in” relation, the relation component in OAR is formulated by $\mathcal{R}_{t}=\{{\bf r}_{1}^{(t)},\cdots,{\bf r}_{K}^{(t)}\}$.
@@ II-C OAR Based Video Generative Reconstruction
At the receiver, generative reconstruction is conducted based on the received OAR sequence combined with reference frames. The overall framework is shown in Fig. 3, which can be divided into four modules: (1) OAR embedding and graph computing, (2) OAR layout generation, (3) optical flow estimation and frame prediction, and (4) image synthesis and fusion. Because the OAR for each video frame is processed by the same process of OAR embedding, graph computing and layout generation, the time superscripts have been omitted in modules (1) and (2) for brevity.
@@ II-C1 OAR Embedding and Graph Computing
For the object “background”, the category is embedded based on the same method, and the angle embedding vector is set to a constant ${\bf 0}$.
A graph convolutional network (GCN, denoted by $\mathcal{G}_{\text{OAR}}$) is utilized with reference to [32] for deep feature extraction of OARs. Specifically, a primary graph is constructed with the concatenation of category and angle embedding vectors ${\bf e}^{(n)}=\text{concat}({\bf e}_{c}^{(n)},{\bf e}_{\theta}^{(n)})$ as nodes and the relation embedding vectors ${\bf e}_{\text{r}}^{(k)}$ as directed edges from $o_{\text{s},k}$ to $o_{\text{o},k}$. Subsequently, the primary graph is processed by a two-layer GCN to obtain a deep feature for each object, denoted by ${\bf f}^{(n)}$.
@@ II-C2 OAR Layout Generation
Temporal dynamics of the motion and scale evolution of objects are captured by the location and size attributes of the OAR. With reference to AG2Video[20], the feature vector of each object is propagated to obtain the layout feature ${\bf L}\in\mathbb{R}^{H\times W\times D}$. The element ${\bf L}_{i,j,d}$ is formulated as:
@@
where ${\bf L}^{(n)}\in\mathbb{R}^{H\times W\times D}$ represents the individual layout of the $n$-th object. Specifically, the deep feature ${\bf f}^{(n)}$ is expanded to the region corresponding to the bounding box ${\bf b}_{n}=(x_{n},y_{n},w_{n},h_{n})$, which is computed by:
@@
One layout is generated form OAR of each video frame, constructing the layout sequence: $[{\bf L}_{1},\cdots,{\bf L}_{T}]$.
II-C3 Optical Flow Estimation and Frame Prediction
With reference frames providing appearance information of objects, primary image prediction is conducted based on the motion information of OAR via optical flow estimation. The $t$-th frame ($t\geq 2$) is predicted based on the previous synthesized frame $\hat{\bf I}_{t-1}$ and adjacent layouts ${\bf L}_{t-1},{\bf L}_{t}$. First, a semantic feature extractor $\mathcal{E}_{\text{s}}$ and a pixel feature extractor $\mathcal{E}_{\text{p}}$ are utilized to obtain deep OAR features of the two frames ${\bf F}_{\text{s}}^{(t)}=\mathcal{E}_{\text{s}}({\bf L}_{t}),{\bf F}_{\text{s}}^{(t-1)}=\mathcal{E}_{\text{s}}({\bf L}_{t-1})$, and image features of the previous frame ${\bf F}_{\text{p}}^{(t-1)}=\mathcal{E}_{\text{p}}(\hat{\bf I}_{t-1})$, respectively. Optical flows are then estimated via an optical flow estimator $\mathcal{F}$ in two stages:
@@ II-C3 Optical Flow Estimation and Frame Prediction
$=\mathcal{F}_{\text{d}}(\mathcal{F}_{\text{e}}({\bf F}_{\text{s}}^{(t)},{\bf F}_{\text{s}}^{(t-1)},{\bf F}_{\text{p}}^{(t-1)}))$
$\in\mathbb{R}^{H\times W\times 2}.$
@@
Figure 4: The framework of OAR-assisted image JSCC. The SNR is assumed to be available at both transmitter and receiver. OAR undergoes lossless transmission, ensuring that both the transmitter and receiver obtain identical OAR features ${\bf F}_{\text{l}}$ through the utilization of OAR feature extraction networks with identical parameters.
@@ II-C4 Image Synthesis and Fusion
While frame prediction enables the reconstitution of object motions, occluded regions and newly emerging objects cannot be resolved. Accordingly, generative reconstruction is employed for detail complementation with reference to vid2vid [19]. Specifically, an image synthesis network $\mathcal{S}$ is implemented to generate the current frame by

### v2

@@ I Introduction
Research into explicit semantic representation and transmission of videos is still limited. Studies such as [16, 17, 18, 19, 20] have explored using semantic segmentation maps and scene graphs for semantic representations, implementing corresponding image and video generation to preserve the videos’ semantic content. Deep learning methods such as object detection[21, 22], classification[23] and relation recognition[24, 25, 26] have been employed for scene semantic parsing and semantic organization. However, these studies did not deeply engage with compression perspectives or data volume reduction. In [27, 28, 8], semantic segmentation maps served as explicit semantic representations for image compression, enhancing semantic fidelity by varying bit-rates across different categories. These efforts treated semantic information as supplementary to traditional image compression, relying still on transmitting numerous pixel and residual features for quality enhancements, which hindered significant bit-rate reductions. [29] took a different approach by encoding and transmitting only semantic segmentation maps and reconstructing images using a diffusion model, catering to extreme channel conditions. Similarly, [11] used sketch representations and generative models. In video conferencing, efficient transmission was achieved using facial keypoints[30]. Nevertheless, these approaches primarily focused on the semantic representation and coding of images or simpler video formats. The challenge remains to explore these techniques for large-scale, time-continuous, and object-consistent video representation to fully address the complexities of large scene videos.
@@
To address the above issues, this paper proposes a novel video transmission system leveraging object-attribute-relation (OAR) representation model of semantic information, as shown in Fig. 1 (b). Specifically, the OAR model characterizes scene semantics by capturing key objects, their attributes, and the relationships within video frames, tailored to the requirements of downstream applications. Benefiting from advancements in existing deep learning algorithms[21, 22, 23, 24], OAR-based semantic information can be efficiently and automatically extracted. Compared with object edges and semantic segmentation maps widely used in previous studies, OAR captures more structured semantic correlations, which are represented by connected graphs; compared with scene graph representation, OAR pays more attention to the enrichment of attributes, which enables better reconstruction of videos. In addition, this paper implements a JSCC pipeline for video transmission by exploiting the advantages of compact and downstream tasks related OAR data. Specifically, this paper implements OAR-assisted JSCC for key frames so that the model can pay more attention to the key objects in the scene. Experimental validation is performed on traffic surveillance videos, and the results indicate that the proposed system can still maintain better foreground reconstruction quality under low bandwidth and SNR conditions.
@@
    Experiments on a traffic surveillance dataset show that the proposed method improves the perceptual quality of reconstructed videos significantly compared to conventional methods such as H.265 and prior work in semantic communication, namely DVST. Specifically, at a CBR of 1/300, the proposed method exhibits an average reduction in LPIPS loss of 0.054 when compared to DVST. Additionally, it facilitates improved performance in object detection tasks, achieving an average increase in mean average precision (mAP) of 0.15 compared to DVST.
@@ II-A Object-Attribute-Relation Formulation and OAR Based Video Communication Framework
An Object-Attribute-Relation (OAR) representation is a structured abstraction of the scene semantics in the form of a graph. For image $\bf I$, the OAR is expressed as a tuple of objects, attributes and relations: ${\bf I}\stackrel{{\scriptstyle\mathcal{E}_{OAR}}}{{\longrightarrow}}\left(\mathcal{O},\mathcal{A},\mathcal{R}\right)$, where $\mathcal{E}_{OAR}$ denotes the mapping from the image to OAR. For a video with $T$ consecutive frames, the OAR is represented as a temporal sequence, $[\left(\mathcal{O}_{1},\mathcal{A}_{1},\mathcal{R}_{1}\right),\cdots,\left(\mathcal{O}_{T},\mathcal{A}_{T},\mathcal{R}_{T}\right)]$.
As shown in the expression, OAR contains three components: ${\mathcal{O}}=[o_{1},\cdots,o_{N}]$, ${\mathcal{A}}=[{\bf a}_{1},\cdots,{\bf a}_{N}]$, and =${\mathcal{R}}=\{r_{1},\cdots,r_{K}\}$, where $N,K$ denote the numbers of objects and relations, respectively. ${\mathcal{O}}$ is the object set with $o_{n}$ denoting a unique ID for each object. ${\mathcal{A}}$ is the attribute set, with ${\bf a}_{n}=\left(a^{(n)}_{1},a^{(n)}_{2},\cdots,a^{(n)}_{A}\right)={\mathcal{E}_{\text{A}}}({\bf I},o_{n})$ denoting $A$ attributes of object $o_{n}$. The relation is denoted by =${\mathcal{R}}=\left\{{\bf r}_{1},\cdots{\bf r}_{R}\right\}=\left\{(o_{\text{s},1},o_{\text{o},1},r_{1}),\cdots,(o_{\text{s},K},o_{\text{o},K},r_{K})\right\}$, $r_{k}=r_{o_{\text{s},k}\rightarrow o_{\text{s},k}}$ representing the relation with subject $o_{\text{s},k}$ and object $o_{\text{o},k}$. An OAR is characterized by a directed graph with objects as nodes and relations as directed edges. The relations and attributes (e.g., category, color, etc.) take values from a deterministic set, i.e.,
@@
OAR is highly interpretable and has a small data volume because only semantics of key objects are characterized. It is especially conducive to task-oriented applications with steady backgrounds, such as surveillance videos. Additionally, OAR provides scalability and can be adapted to complex scenarios through the extension of attributes and relations. Based on the above modeling, this article proposes an OAR-based video transmission framework, as presented in Fig. 2.
In the proposed framework, $T$ consecutive frames are encapsulated as one group-of-pictures (GoP). The first frame of each GoP (known as the reference frame or key frame) is intra-coded and transmitted, providing contour and texture information about the background and key objects. The semantics of all frames in a GoP are represented as OARs and transmitted after entropy coding, channel coding and modulation. OAR sequences are decoded after demodulation at the receiver, and generative reconstruction is carried out based on the reference frame incorporated with OAR sequences. Without loss of generality, each GOP is independently processed. The reference frames are denoted by ${\bf I}_{1}$, and subsequent $T-1$ frames are denoted by ${\bf I}_{t},t=2,\cdots,T$. For objects contained in reference frames, faithful reconstruction is realized based on the motion information in OAR. Newly appeared objects are generated based on the prior knowledge of the trained models, resulting in possible bias in pixel-level details such as colors and textures. This article focuses on task-oriented video transmission, so deviations in detailed textures are acceptable under the premise of semantic fidelity.
II-B OAR Extraction and Representation
OAR representation is the critical component of the proposed video transmission system, providing semantics with low bit-rate representation and enabling efficient video transmission.Semantic concepts require specifications depending on different scenarios and tasks. Consider a surveillance video transmission system oriented toward tasks such as object detection and anomaly monitoring. Vehicles in the foreground are treated as the objects in OAR. The attributes consist of the position ${\bf p}=(x,y)$, size ${\bf s}=(w,h)$, angle $\theta$, and category $c$, denoted by ${\bf a}=(a_{1},a_{2},a_{3},a_{4})=({\bf p},{\bf s},\theta,c)$. Relations involve mainly “occlusions” between foreground objects and occlusions of the foreground by the background. The trivial relation of objects “in” the background is also included. Consequently, for an image of size $W\times H$, the attributes and relations take the value space of:
@@ II-B OAR Extraction and Representation
To realize automated OAR extraction, multi-object tracking (MOT) is utilized, followed by several attribute recognition models. The overall process consists of three parts: multi-object tracking, attribute recognition, and relation identification.
First, multi-object tracking. Object detection is first conducted via YOLOv9[34] to obtain the categories $c_{n}^{(t)}$ and bounding boxes ${\bf b}_{n}^{(t)}=(x_{n}^{(t)},y_{n}^{(t)},w_{n}^{(t)},h_{n}^{(t)})$, where $n=1,\cdots,N_{t}$. To match the identical objects in consecutive frames, a unique ID $o_{n}^{(t)}$ is assigned to each object via the SORT algorithm[35]. IDs for all objects constitute the object component ${\mathcal{O}_{t}}=[o_{1}^{(t)},\cdots,o_{N_{t}}^{(t)}]$ in OAR.
@@
Third, relation identification. Object pairs with intersecting bounding boxes are examined and the occlusion relations are determined by comparing the appearance of the overlapping regions with the respective object features. In addition, occlusions between objects and background elements are also considered. Regions located at the spatial forefront are manually labeled. Objects overlapping with the labeled area are considered to be occluded by the background. By incorporating the identified “occlusion” and the trivial “in” relation, the relation component in OAR is formulated as $\mathcal{R}_{t}=\{{\bf r}_{1}^{(t)},\cdots,{\bf r}_{K}^{(t)}\}$.
@@ II-C OAR Based Video Generative Reconstruction
At the receiver, generative reconstruction is conducted based on the received OAR sequence combined with reference frames. The overall framework is shown in Fig. 3, which can be organized into four modules: (1) OAR embedding and graph computing, (2) OAR layout generation, (3) optical flow estimation and frame deformation, and (4) image synthesis and fusion. Because the OAR for each video frame is processed via the same process of OAR embedding, graph computing and layout generation, the time superscripts are omitted in modules (1) and (2) for brevity.
@@ II-C1 OAR Embedding and Graph Computing
For the object “background”, the category is embedded via the same method, and the angle embedding vector is set to a constant ${\bf 0}$.
Although OAR realizes the extraction of attributes and relations between objects in the scene, there are still close connections and interactions between objects. This cross attention across space and individuals inspires us to perform propagation and modulation of features across objects after embedding, and the relations are an important basis for this process. Considering graph convolutional networks (GCNs) have been utilized to extract high-level semantic features from graph-structured data and have shown excellent performance on classification tasks [38, 39] and generative tasks [40, 41] of graph data, a GCN (denoted by $\mathcal{G}_{\text{OAR}}$) is utilized with reference to [38] for deep feature extraction of OARs. Specifically, a primary graph is constructed with the concatenation of category and angle embedding vectors ${\bf e}^{(n)}=\text{concat}({\bf e}_{c}^{(n)},{\bf e}_{\theta}^{(n)})$ as nodes and the relation embedding vectors ${\bf e}_{\text{r}}^{(k)}$ as directed edges from $o_{\text{s},k}$ to $o_{\text{o},k}$. The primary graph is subsequently processed by a two-layer GCN to obtain a deep feature for each object, denoted by ${\bf f}^{(n)}$. Multilayer GCNs further expand the scope of attention so that the obtained features can be affected by the associated multiple objects simultaneously.
@@ II-C2 OAR Layout Generation
The temporal dynamics of the motion and scale evolution of objects are captured by the location and size attributes of the OAR. With reference to AG2Video[20], the feature vector of each object is propagated to obtain the layout feature ${\bf L}\in\mathbb{R}^{H\times W\times D}$. The element ${\bf L}_{i,j,d}$ is formulated as:
@@
where ${\bf L}^{(n)}\in\mathbb{R}^{H\times W\times D}$ represents the individual layout of the $n$-th object. Specifically, the deep feature ${\bf f}^{(n)}$ is expanded to the region corresponding to the bounding box ${\bf b}_{n}=(x_{n},y_{n},w_{n},h_{n})$, which is computed as follows:
@@
One layout is generated from the OAR of each frame, constructing the layout sequence: $[{\bf L}_{1},\cdots,{\bf L}_{T}]$.
II-C3 Optical Flow Estimation and Frame Deformation
With reference frames providing appearance information of objects, optical flows are estimated based on the motion information of OAR and frame deformation is conducted. The $t$-th frame ($t\geq 2$) is predicted based on the previously synthesized frame $\hat{\bf I}_{t-1}$ and adjacent layouts ${\bf L}_{t-1},{\bf L}_{t}$. First, a semantic feature extractor $\mathcal{E}_{\text{s}}$ and a pixel feature extractor $\mathcal{E}_{\text{p}}$ are utilized to obtain deep OAR features of the two frames ${\bf F}_{\text{s}}^{(t)}=\mathcal{E}_{\text{s}}({\bf L}_{t}),{\bf F}_{\text{s}}^{(t-1)}=\mathcal{E}_{\text{s}}({\bf L}_{t-1})$, and image features of the previous frame ${\bf F}_{\text{p}}^{(t-1)}=\mathcal{E}_{\text{p}}(\hat{\bf I}_{t-1})$, respectively. Optical flows $\hat{\bf F}_{\text{m}}^{(t)}\in\mathbb{R}^{H\times W\times 2}$ are then estimated via an optical flow estimator $\mathcal{F}$ in two stages:
@@ II-C3 Optical Flow Estimation and Frame Deformation
$=\mathcal{F}_{\text{d}}(\mathcal{F}_{\text{e}}({\bf F}_{\text{s}}^{(t)},{\bf F}_{\text{s}}^{(t-1)},{\bf F}_{\text{p}}^{(t-1)})).$
@@
Figure 4: The framework of OAR-assisted image JSCC. The SNR is assumed to be available at both the transmitter and the receiver. OAR undergoes lossless transmission, ensuring that both the transmitter and receiver obtain identical OAR features ${\bf F}_{\text{l}}$ through the utilization of OAR feature extraction networks with identical parameters.
@@ II-C4 Image Synthesis and Fusion
While frame deformation enables motion reconstitution, occluded regions and newly emerging objects cannot be resolved. Accordingly, generative reconstruction is employed for detail complementation with reference to vid2vid [19]. Specifically, an image synthesis network $\mathcal{S}$ is implemented to generate the current frame by
@@
Denote the parameters of the network to be optimized as $F=\{{\bf W}_{\text{c}},{\bf W}_{\theta},{\bf W}_{\text{r}},\mathcal{G}_{\text{OAR}},\mathcal{E}_{\text{s}},\mathcal{E}_{\text{p}},\mathcal{S},\mathcal{F}_{\text{e}},\mathcal{F}_{\text{d}},\mathcal{W}\}$, and the modules are jointly trained following [19] by solving:
$\min_{F}\left(\max_{D_{I}}\mathcal{L}_{I}(F,D_{I})+\max_{D_{V}}\mathcal{L}_{V}(F,D_{V})\right)+\lambda_{W}\mathcal{L}_{W}(F),$ (9)
where subscripts $I$ and $V$ denote the optimization of single and consecutive video frames, respectively. $D_{I/V}$ denotes the discriminator, and $\mathcal{L}_{I/V}$ denotes the weighted summation of the VGG loss and the adversarial loss. $\mathcal{L}_{W}$ denotes optical flow loss, which guides the training process of the optical flow estimation module $\mathcal{F}_{\text{e}}\cdot\mathcal{F}_{\text{d}}$ in terms of the accuracy of the optical flow $\hat{\bf F}_{\text{m}}$ and the quality of the deformed image $\check{\bf I}_{t}$.